OpenStudy (anonymous):

p= 2l+2w; L=17, w=56 p=

OpenStudy (diyadiya):

p=2(17)+2(56)=34+112=146

OpenStudy (anonymous):

146 = 2(17) + 2(56)

OpenStudy (anonymous):

Nice job Diya

OpenStudy (diyadiya):

Thaks:)